OpenStudy (anonymous):

Find the triple integral of (1-z^2)dxdydz of the pyramid. D = [0,1]x[0,1]x[0,1]

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{1}\int\limits_{0}^{1} (1-z^{2})dx dy dz\]

OpenStudy (anonymous):

integrate wrt x first, as it is dx dy dz

OpenStudy (anonymous):

did you get 2/3?

OpenStudy (anonymous):

well, not really. I can't get the book's answer of 3/10

OpenStudy (anonymous):

with respect to x, i would get x-xz^2 evaluated from 0 to 1. which brings me back to 1-z^2 dydz

OpenStudy (anonymous):

are you sure the question is given as (1-z^2) dx dy dz or do we have (1-z^2) dz dy dx

OpenStudy (anonymous):

i am sure

OpenStudy (anonymous):

wait a minute, this is pyramid, okay so here is what you need to do (1-z^2) is bounded by 0 and 1 , now make a project of this 3D figure in 2D ,

OpenStudy (anonymous):

the whole question is (1-z^2)dxdydz; W is a pyramid with the top vertex of (0,0,1) and base vertices of (0,0,0), (1,0,0), (0,1,0) and (1,1,0)

OpenStudy (anonymous):

and bound of y and x need to be different

OpenStudy (anonymous):

oh pellet, change of order integration?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the reason, you dxdydz, so we need either dzdydx or dzdxdy, hope that helps, i am sure youwill have answer 3/10

OpenStudy (anonymous):

how would that change the bounds of x and y? would they not still be 0 to 1?

OpenStudy (anonymous):



OpenStudy (anonymous):

figure not to scale

OpenStudy (anonymous):

so z varies from origin to the top most point, and then you can have either projection in XY plane OR CHOOSE OTHER ALTERNATIVEs

OpenStudy (anonymous):

ok. thank you